Let and consider V={x∈R^2 | Ax=5x}. Prove that V is a subspace of R^2, find a...

Let Chv3POTZ2F8AAAAASUVORK5CYII= and consider V={xR^2 | Ax=5x}. Prove that V is a subspace of R^2, find a basis for V, and determine its dimension.
Answer #1

* Given Itat A-[11] V = { XEvR) Ax =5x}. now, we have to show that visa subspace of 1R². so AX = 54, AY=57. A(x-y) = AX-AY =AX = 5x [97] [%] =5[3] [924) -5[*] = [0] (1928-59) - [O] » [ ***] = [] *1 *[]+4[*] = [] · Basis = {[7] [24]} dimension = num